# Quantum tunneling from excited states: Recovering imaginary-time instantons from a real-time analysis

### v1

@@ IV Regularized real-time dynamics
A similarly simple interpretation can be obtained for the case of a small but finite $\epsilon$, for which the path can be decomposed into a real and imaginary part, $\bar{x}(t)=\bar{x}_{\rm Re}(t)+\bar{x}_{\rm Im}(t)$. Their corresponding equations of motions are given by
@@ V.1 Periodic steadyon solutions
The transition of the particle from one basin to the other amounts to evolution through one-half period of the periodic motion along the imaginary-time direction. Together with the crossing condition, this implies that a transition from $x_{i}$ to $x_{s}$ after the time $T$ would require
@@
To address this issue, we first recall that the impossibility to perform the stationary-phase approximation using the solutions of Eq. (27) for some physical time $T$ does not have an immediate physical meaning—the decay rate is still well-defined, we just cannot compute it in the way we were hoping for. However, as we will see in the next subsection, knowledge of the decay rate on a countable set of times is sufficient to deduce the physical decay rate in the limit of interest. Starting from this premise, we can focus our attention on those discrete times for which it is at least in principle possible to satisfy the boundary conditions.
@@
We may now recall that one of the main motivations for our analysis was the simple observation that the usual interpretation of imaginary time as Wick-rotated physical time seemed to suggest that an instanton can only exist at one finite physical time, of order of the system’s natural time-scale. At first glance, this appears to be true even in our more careful treatment, as there is still only one unique value for $t_{s}$ for any given $\epsilon$. This apparent problem can be resolved thanks to two simple observations. First, we note that the relation between $x_{i}$, $t_{s}$, and $\epsilon$ necessary to (approximately) satisfy the boundary condition can be understood equivalently as determining $\epsilon$ in terms of the physical parameters $x_{i}$ and $t_{s}$, with $\epsilon\sim\Delta\tau_{I}/(2t_{s})$. Second, we recall that in our picture, $\epsilon$ is not an entirely free parameter, but corresponds to a regularization of the Hamiltonian, and hence needs to be taken to 00 at the end of a given calculation. These two points suggest that for any suitable time $t_{s}$ of interest, we may choose the “right” regularization parameter $\epsilon$ to ensure the existence of an approximate stationary phase. Inevitably, any such choice implies a finite error. However, as we consider larger and larger physical times (in discrete steps), we also lower the value of $\epsilon$ required, implying an increasingly better approximation. In the limit $T=t_{s}\to\infty$, this automatically yields $\epsilon\to 0$.
@@ V.2 Non-periodic steadyon solutions
Returning to our interpretation of the steadyon as two interacting degrees of freedom, we can immediately understand the effect of this initial velocity. For the periodic solution, we had found that the tunneling occurs as the oscillations of $\bar{x}_{\rm Re}(t)$ and $\bar{x}_{\rm Im}(t)$ drive each other until $\bar{x}_{\rm Re}(t)$ crosses from the false-vacuum basin $\Omega_{\rm FV}$ into the other well $\Omega_{R}$. Upon giving $\bar{x}_{\rm Im}(t)$ an initial velocity, this process will unfold faster for a given value of $\epsilon$, also leading to a smaller $\Delta\tau_{I}$. See Fig. 6.
@@
The realization of these boundary conditions can be achieved in a similar way as for the periodic case. For general physical times, no exact solution exists. There exists, however, a countable number of physical times for which $\bar{x}_{\rm Im}(t)$ vanishes and $\bar{x}_{\rm Re}(t)$ takes a value close to the emergence point $x_{s}$, which agrees with the value of the instanton at $\tau=\epsilon\cdot T$. This offset is controlled by $\epsilon$, and therefore vanishes in the limit $\epsilon\to 0$, $T\to\infty$, consistent with Eq. (32).
@@ V.3 Real-time action and instanton limit
In order to illustrate this behavior, we calculate the imaginary part of the steadyon action for multiple values of $\epsilon$ by simply inserting the corresponding solutions into the regularized action of Eq. (22) with the appropriate $t_{s}(\epsilon)$ for both the periodic steadyon as well as the steadyon with nonzero initial velocity, which we will later find to dominate the tunneling rate. And indeed, taking the limit in the suggested way leads to a clear convergence. Crucially, we find that the limiting value is given precisely by the Euclidean action of the corresponding instanton in the imaginary-time picture. See Figs. 8 and 8.
@@
$t_{s}$$t$$\tau$Im$\left(S\right)=0$$t_{s}\cdot\epsilon$$O(\epsilon)$$\Delta\tau_{I}/2$$\epsilon$
@@
However, unlike for the periodic steadyon, we find that the analytic continuation onto the real-time axis is not a strictly real function linked to classical behavior, as shown in Fig. 11. This, in turn, yields a complex Lagrange density, see Fig. 11. First we note that the double-periodicity of our more general solutions of Eq. (27) manifests, again, on the real-time axis. Next we observe that within each period of length $\Delta t_{R}$, the Lagrange function is asymmetric, such that its net contribution to the action per interval of this length vanishes. Following our discussion around Eq. (31), the boundary conditions imply that, in the limit $\epsilon\to\infty$, $t_{s}\to N\cdot\Delta t_{R}$, with some positive integer $N$. Thus, although the Lagrange function on the real-time axis is not manifestly real itself, its overall contribution to ${\rm Im}(S[\bar{x}])$ vanishes in the appropriate limit. As the contribution from the imaginary-time axis reproduces, by construction, that of the instanton, we recover Eq. (33).
@@ VI Normalization factor
Having understood the numerator of the decay rate, it remains to evaluate the denominator, i.e., the normalization factor $Z=P_{\Omega}$. Once again, we wish to do so by means of the stationary-phase approximation, which requires us to match the solutions of Eq. (27) with the boundary conditions of Eq. (26). Before doing so for the real-time case, it is helpful for the interpretation of our results to consider the imaginary-time picture, which we will ultimately recover in the relevant limits.
@@ VI.1 Pulls
In the limit $x_{i}\to x_{\rm FV}$, the saddle point of the normalization factor is simply given by $\bar{z}_{I}(\tau)=x_{\rm FV}$. We furthermore anticipate that the saddle-point solution will be defined on an imaginary-time interval of the same length as the instanton, $\Delta\tau_{I}$.
This suggests that we consider solutions that begin at $x_{i}$ with an initial velocity that is just large enough for the particle to reach some turning point $x_{p}$ at $\tau=\Delta\tau_{I}/2$ and return to $x_{i}$ at $\tau=\Delta\tau_{I}$, as in Fig. 13. We refer to such solutions as “pulls”: they begin at $x_{i}$ and are “pulled back” to some point $x_{v}$ that is also within the false-vacuum basin $\Omega_{\rm FV}$. We find that, for any pair $\Delta\tau_{I}$ and $x_{i}$, these solutions minimize the Euclidean action, that is, they give rise to the dominant contribution to the normalization factor. This allows us to evaluate the integral over the steadyon’s end point $x_{v}$ in Eq. (18) using a saddle-point approximation around the value of $x_{v}$ corresponding to this behavior.
@@ VI.3 Real-time action and pull limit
Figure 16: Left panel: The Euclidean action of the dominant instanton as a function of the initial positon $x_{i}$. Middle panel: The Euclidean action of the pull corresponding to the dominant instanton as a function of the initial positon $x_{i}$. Right panel: The leading order exponent of the decay rate as a function of the initial positon $x_{i}$.
@@ VI.4 Effect of the pull on the leading-order decay rate
We can distinguish two regimes for the pull, based on the value of $x_{p}$. First, if $x_{i}$ is sufficiently close to $x_{\rm FV}$ and $\Delta\tau_{I}$ is large enough, the Euclidean action is minimized for values of $x_{p}$ close to $x_{\rm FV}$. This implies that, for most of the time interval $\Delta\tau_{I}$, the particle hovers near $x_{\rm FV}$, corresponding to a near-vanishing Euclidean Lagrangian. In order to get to this point, the particle needs a large enough initial velocity, giving rise to a contribution to the Euclidean action from the kinetic term. If, on the other hand, $x_{i}$ is further away from $x_{\rm FV}$ and $\Delta\tau_{I}$ is small enough, the reduction in Euclidean action achieved by approaching $x_{p}$ cannot compensate for this additional contribution. In the most extreme case of a near-vanishing $\Delta\tau_{I}$, this can lead to the pull essentially approaching a constant function with $x_{p}\simeq x_{i}$, with a vanishing Euclidean action.
@@ VII Conclusion
We have derived a complete path-integral expression for the tunneling rate out of an excited state from first principles. On a technical level, our most important result is a strategy for the evaluation of path integrals in real time. To avoid the complications linked to a Wick rotation by a finite angle, we regularize the path integral by introducing a small imaginary part for the Hamiltonian, controlled by a regularization parameter $\epsilon$. This procedure is well-defined in both quantum mechanics and quantum field theory, and has also been shown to produce meaningful results in curved backgrounds and even for systems with an explicit time-dependence [15].
For systems without an explicit time-dependence, this step is equivalent to an infinitesimal Wick rotation onto a complex-time contour $\gamma_{\epsilon}$. This regularization allows us to identify a countable number of points in the $(\epsilon,T)$-plane for which the path integral can be evaluated using the stationary-phase method. We refer to the (approximate) stationary-phase solutions as steadyons. Most important, these points approach the real-time axis in the limit of a vanishing regulator. Assuming continuity of the decay rate, the results obtained using the steadyons then converge to the physical decay rate in the combined limit $\epsilon\to 0,\,T\to\infty$. We find that, for systems without an explicit time-dependence, this reproduces the familiar result obtained in the usual imaginary-time picture. This requires the steadyon’s analytic continuation to be doubly-periodic and regular between the real-time axis and $\gamma_{\epsilon}$. This can be expected for a wide class of physical systems and is easily confirmed for the imaginary-time instanton, allowing for a straightforward generalization of our result.
@@
While motivated by the need to realize the boundary conditions describing quantum tunneling, our method can also be used to evaluate more general path integrals with non-trivial boundary conditions. This includes, in particular, the normalization factor of the tunneling rate. In the real-time picture, the latter is dominated by solutions which we called drags, which we have shown to be linked to imaginary-time solutions called pulls in a similar way as the steadyons are related to instantons. The “drag” contributions stem from summing over paths that remain within the false-vacuum basin $\Omega_{\rm FV}$.
@@ Acknowledgements.
We thank Wen-Yuan Ai, Morgane König, and Matthew D. Schwartz for valuable discussions. TS’s contributions to this work were made possible by the Walter Benjamin Programme of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 512630918. Portions of this work were conducted in MIT’s Center for Theoretical Physics and partially supported by the U.S. Department of Energy under Contract No. DE-SC0012567. This project was also supported in part by the Black Hole Initiative at Harvard University, with support from the Gordon and Betty Moore Foundation and the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of these Foundations.

### v2

@@ IV Regularized real-time dynamics
A similarly simple interpretation can be obtained for the case of a small but finite $\epsilon$, for which the path can be decomposed into a real and imaginary part, $\bar{x}(t)=\bar{x}_{\rm Re}(t)+i\bar{x}_{\rm Im}(t)$. Their corresponding equations of motions are given by
@@ V.1 Periodic steadyon solutions
The transition of the particle from one basin to the other amounts to an evolution through one-half period of the periodic motion along the imaginary-time direction. Together with the crossing condition, this implies that a transition from $x_{i}$ to $x_{s}$ after the time $T$ would require
@@
This is, however, not surprising, but rather a manifestation of the physical motivation for our ansatz in Eq. (7). Recall that we had argued that the loss of probability within $\Omega_{-}$ should happen predominantly at times for which the point particle is furthest up the potential barrier, i.e., $T=N\cdot\Delta t_{R}$. And indeed, this is precisely the behavior shown by the steadyon.
@@
We may now recall that one of the main motivations for our analysis was the simple observation that the usual interpretation of imaginary time as Wick-rotated physical time seemed to suggest that an instanton can only exist at one finite physical time, of order of the system’s natural time-scale. At first glance, this appears to be true even in our more careful treatment, as there is still only one unique value for $t_{s}$ for any given $\epsilon$. This apparent problem can be resolved thanks to two simple observations. First, we note that the relation between $x_{i}$, $t_{s}$, and $\epsilon$ necessary to (approximately) satisfy the boundary condition can be understood equivalently as determining $\epsilon$ in terms of the physical parameters $x_{i}$ and $t_{s}$, with $\epsilon\sim\Delta\tau_{I}/(2t_{s})$. Second, we recall that in our picture, $\epsilon$ is not an entirely free parameter, but corresponds to a regularization of the Hamiltonian, and hence needs to be taken to 0 at the end of a given calculation. These two points suggest that for any suitable time $t_{s}$ of interest, we may choose the “right” regularization parameter $\epsilon$ to ensure the existence of an approximate stationary phase. Inevitably, any such choice implies a finite error. However, as we consider larger and larger physical times (in discrete steps), we also lower the value of $\epsilon$ required, implying an increasingly better approximation. In the limit $T=t_{s}\to\infty$, this automatically yields $\epsilon\to 0$. See Fig. 6.
@@ V.2 Non-periodic steadyon solutions
Returning to our interpretation of the steadyon as two interacting degrees of freedom, we can immediately understand the effect of this initial velocity. For the periodic solution, we had found that the tunneling occurs as the oscillations of $\bar{x}_{\rm Re}(t)$ and $\bar{x}_{\rm Im}(t)$ drive each other until $\bar{x}_{\rm Re}(t)$ crosses from the false-vacuum basin $\Omega_{-}$ into the other well $\Omega_{+}$. Upon giving $\bar{x}_{\rm Im}(t)$ an initial velocity, this process will unfold faster for a given value of $\epsilon$, also leading to a smaller $\Delta\tau_{I}$. See Fig. 7.
@@
The realization of these boundary conditions can be achieved in a similar way as for the periodic case. For general physical times, no exact solution exists, in agreement with our initial ansatz. Changing the initial momentum of the instanton also changes the length of the interval $\Delta t_{R}$. This manifests in our calculation through the condition that the steadyon only exists for precisely these times. We can also understand this behavior more generally in terms of the usual interpretation of the path integral as a superposition of different paths. The position eigenstate has no definite momentum, such that the paths within the path integral setting out from that position contain all possible momenta. Thus, the simple picture given in Fig. 2 can be understood as describing just one of these paths, for which the particle’s momentum at $x=x_{i}$ vanishes.
@@ V.3 Real-time action and instanton limit
In order to illustrate this behavior, we calculate the imaginary part of the steadyon action for multiple values of $\epsilon$ by simply inserting the corresponding solutions into the regularized action of Eq. (22) with the appropriate $t_{s}(\epsilon)$ for both the periodic steadyon as well as the steadyon with nonzero initial velocity, which we will later find to dominate the tunneling rate. And indeed, taking the limit in the suggested way leads to a clear convergence. Crucially, we find that the limiting value is given precisely by the Euclidean action of the corresponding instanton in the imaginary-time picture. See Figs. 9 and 9.
@@
[…]
@@
However, unlike for the periodic steadyon, we find that the analytic continuation onto the real-time axis is not a strictly real function linked to classical behavior, as shown in Fig. 12. This, in turn, yields a complex Lagrange density, see Fig. 12. First we note that the double-periodicity of our more general solutions of Eq. (27) manifests, again, on the real-time axis. Next we observe that within each period of length $\Delta t_{R}$, the Lagrange function is asymmetric, such that its net contribution to the action per interval of this length vanishes. Following our discussion around Eq. (31), the boundary conditions imply that, in the limit $\epsilon\to\infty$, $t_{s}\to N\cdot\Delta t_{R}$, with some positive integer $N$. Thus, although the Lagrange function on the real-time axis is not manifestly real itself, its overall contribution to ${\rm Im}(S[\bar{x}])$ vanishes in the appropriate limit. As the contribution from the imaginary-time axis reproduces, by construction, that of the instanton, we recover Eq. (33). Crucially, this discussion is independent of our choice of potential, and in particular does not require the potential to have degenerate vacua.
@@ VI Normalization factor
Having understood the numerator of the decay rate, it remains to evaluate the denominator, i.e., the normalization factor $Z=P_{\Omega_{-}}$. Once again, we wish to do so by means of the stationary-phase approximation, which requires us to match the solutions of Eq. (27) with the boundary conditions of Eq. (26). Before doing so for the real-time case, it is helpful for the interpretation of our results to consider the imaginary-time picture, which we will ultimately recover in the relevant limits.
@@ VI.1 Pulls
In the limit $x_{i}\to x_{-}$, the saddle point of the normalization factor is simply given by $\bar{z}_{I}(\tau)=x_{-}$. We furthermore anticipate that the saddle-point solution will be defined on an imaginary-time interval of the same length as the instanton, $\Delta\tau_{I}$.
This suggests that we consider solutions that begin at $x_{i}$ with an initial velocity that is just large enough for the particle to reach some turning point $x_{p}$ at $\tau=\Delta\tau_{I}/2$ and return to $x_{i}$ at $\tau=\Delta\tau_{I}$, as in Fig. 14. We refer to such solutions as “pulls”: they begin at $x_{i}$ and are “pulled back” to some point $x_{v}$ that is also within the false-vacuum basin $\Omega_{-}$. We find that, for any pair $\Delta\tau_{I}$ and $x_{i}$, these solutions minimize the Euclidean action, that is, they give rise to the dominant contribution to the normalization factor. This allows us to evaluate the integral over the steadyon’s end point $x_{v}$ in Eq. (18) using a saddle-point approximation around the value of $x_{v}$ corresponding to this behavior.
@@ VI.3 Real-time action and pull limit
Figure 17: Left panel: The Euclidean action of the dominant instanton as a function of the initial positon $x_{i}$. Middle panel: The Euclidean action of the pull corresponding to the dominant instanton as a function of the initial positon $x_{i}$. Right panel: The leading order exponent of the decay rate as a function of the initial positon $x_{i}$. The initial position $x_{1}=-1$ corresponds to tunneling out of the false vacuum. In this limit, the Euclidean action of the pull $S_{E}[\bar{z}]$ vanishes and that of the instanton, $S_{E}[\bar{x}]$, reproduces that of the familiar bounce.
@@ VI.4 Effect of the pull on the leading-order decay rate
We can distinguish two regimes for the pull, based on the value of $x_{p}$. First, if $x_{i}$ is sufficiently close to $x_{-}$ and $\Delta\tau_{I}$ is large enough, the Euclidean action is minimized for values of $x_{p}$ close to $x_{-}$. This implies that, for most of the time interval $\Delta\tau_{I}$, the particle hovers near $x_{-}$, corresponding to a near-vanishing Euclidean Lagrangian. In order to get to this point, the particle needs a large enough initial velocity, giving rise to a contribution to the Euclidean action from the kinetic term. If, on the other hand, $x_{i}$ is further away from $x_{-}$ and $\Delta\tau_{I}$ is small enough, the reduction in Euclidean action achieved by approaching $x_{p}$ cannot compensate for this additional contribution. In the most extreme case of a near-vanishing $\Delta\tau_{I}$, this can lead to the pull essentially approaching a constant function with $x_{p}\simeq x_{i}$, with a vanishing Euclidean action.
@@ VII Conclusion
We have studied the calculation of tunneling rates from first principles using real-time techniques. On a technical level, our most important result is a strategy for the evaluation of the path integrals expected to appear in such calculations. To avoid the complications linked to a Wick rotation by a finite angle, we regularize the path integral by introducing a small imaginary part for the Hamiltonian, controlled by a regularization parameter $\epsilon$. This procedure is well-defined in both quantum mechanics and quantum field theory, and has also been shown to produce meaningful results in curved backgrounds and even for systems with an explicit time-dependence [15].
For systems without an explicit time-dependence, this step is equivalent to an infinitesimal Wick rotation onto a complex-time contour $\gamma_{\epsilon}$. We find that for any physical time $T$ subject to the usual assumptions, there exist choices of $\epsilon$ representing a one-parameter family of approximate saddle points. Crucially, in the limit $T\to\infty$, the existence of these saddle points requires $\epsilon\to 0$, consistent with the role of $\epsilon$ as a regulator. For our (unphysical) choice of initial state, we find an additional condition on the physical times for which these saddle points exist, in agreement with the interpretation of the state.
We find that, for systems without an explicit time-dependence, the relevant limits reproduce the familiar result obtained in the usual imaginary-time picture. This requires the steadyon’s analytic continuation to be doubly-periodic and regular between the real-time axis and $\gamma_{\epsilon}$. This can be expected for a wide class of physical systems and is easily confirmed for the imaginary-time instanton, allowing for a straightforward generalization of our result.
Our discussion focuses primarily on tunneling within a given potential for arbitrary initial positions. One could deploy the formalism we develop to further study how the tunneling rate would change as one modifies the potential as well. Our analysis strongly suggests that the effects of varying the height of the potential barrier would affect the final results in a way similar to what one would expect from the WKB approximation. Overall, we have developed a general formalism with which one could vary both the initial conditions and the particular shape of the potential.
@@
While motivated by the need to realize the boundary conditions describing quantum tunneling, our method can also be used to evaluate more general path integrals with non-trivial boundary conditions. This includes, in particular, the normalization factor of the tunneling rate. In the real-time picture, the latter is dominated by solutions which we called drags, which we have shown to be linked to imaginary-time solutions called pulls in a similar way as the steadyons are related to instantons. The “drag” contributions stem from summing over paths that remain within the false-vacuum basin $\Omega_{-}$.
@@ Acknowledgements.
We thank Wen-Yuan Ai, Morgane König, Bruno Scheihing and Matthew D. Schwartz for valuable discussions, as well as Nils Wagner for pointing out several possibly unclear formulations in an earlier version. TS’s contributions to this work were made possible by the Walter Benjamin Programme of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 512630918. Portions of this work were conducted in MIT’s Center for Theoretical Physics and partially supported by the U.S. Department of Energy under Contract No. DE-SC0012567. This project was also supported in part by the Black Hole Initiative at Harvard University, with support from the Gordon and Betty Moore Foundation and the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of these Foundations.